OpenStudy (anonymous):

2/4 (x- 12)=-4 Is the answer 8?

OpenStudy (mew55):

the answer is 4

OpenStudy (anonymous):

What did I do wrong? :/

OpenStudy (espex):

How did you go about solving it?

OpenStudy (anonymous):

2/4x-8=-4 then i did: 2/4/ 2/4 and -4/24 and divided.

OpenStudy (mew55):

2/4=1/2 by simplifying. 1/2*12=6 6+-4=2 2/1/2=2*2/1=4

OpenStudy (espex):

2/4 * 12 is not -8

OpenStudy (espex):

\[2/4 (x- 12) \rightarrow \frac{2}{4}x - \frac{2*12}{4} \rightarrow \frac{2}{4} - \frac{24}{4}\]

OpenStudy (espex):

Reducing your 24/4 is where you went wrong.

OpenStudy (mew55):

but 2/4 is 1/2 and 1/2*12=6

OpenStudy (espex):

@mew55 Your are correct, @Ano23oe reduced 2/4 * -12 down to -8, that is what I was referencing.

OpenStudy (anonymous):

Wait so then how does 24/4 reduce down to 4?

OpenStudy (espex):

@Ano23oe it doesn't. It reduces down to -6, so you end up with 2/4x-6=-4 and when you solve for x you get 4.

OpenStudy (mew55):

when u get 2/4, it is easier for u to simplify to 1/2

OpenStudy (anonymous):

Ok... This is kind of hard, but I'm going to practice it more thanks! Do you guys mind making up some problems for me to solve?

OpenStudy (espex):

1.) 7 + 3x - 2x = 6 2.) 7 + 7x + 4x = 62 3.) 7x - 1 + 4x = -1 4.) - 47 = x + 6x + 9 5.) 2 + 6x - 4x = 20 6.) 3x + 9 + 5x = 49 7.) 44 = 6x - x - 6 8.) - 6x + 4 - 2x = - 60 9.) x + 2 - 5x = - 30 10.) 6x - 3 - 2x = 29 11.) 4 + x + 7x = - 44 12.) 62 = 7x + 2x - 1 13.) x - 1 + x = 15 14.) 102 = - 7x - 3x + 2 15.) - 4x - 6 + 7x = 9 16.) 5 = 5x + 3x - 3 17.) 40 = 4x + 3x - 2 18.) 2 + 4x + 4x = 98 19.) - 4x + 2 - x = 7 20.) 2x + 8 + 5x = - 55 21.) - 4x - 3 - 5x = - 3 22.) - 4 + x + 4x = 1 23.) - 6x + 4 + 2x = - 12 24.) 0 = x + x + 10 25.) 7x - 4 + 7x = - 32 26.) 7( - 2x + 7) = - 7 27.) 2(6x + 4) = 152 28.) 4(2x + 10) = 112 29.) - 7(3 - 6x) = 483 30.) 5(2 + 7x) = 115 31.) 3(5x - 5) = 60 32.) 3(7x + 8) = 192 33.) - 2( - 2 + 2x) = - 16 34.) 2( - 4x + 4) = - 48 35.) - 3(1x + 7) = - 24 36.) 7( - 2 + x) = - 42 37.) 5( - 8 + 6x) = 20 38.) 2(5 - 6x) = 10 39.) 3(5x + 6) = 18 40.) - 6( - 9 + 5x) = - 66 41.) 2(-1x - 3) = - 24 42.) 4(1 + 7x) = 116 43.) - 7(7 - 2x) = - 91 44.) 7( - 4 - x) = - 84 45.) - 2(5x + 10) = - 90 46.) 3( - 4x - 5) = - 27 47.) 5(6 - 4x) = - 130 48.) 6(1x + 6) = 66 49.) 3(8 + 5x) = 99 50.) - 2( - 7x - 5) = - 88

OpenStudy (espex):

All you could ever want. http://www.math.com/students/worksheet/algebra_sp.htm